# Decentralized Proactive Model Offloading and Resource Allocation for Split and Federated Learning

### v1

@@ 3.4 Regression-based modeling methodology
TABLE I: The Proposed Regression-based Models
4 Problem formulation
In resource-constrained heterogeneous edge environment, considering the heterogeneity of IoT devices’ resource capacities and their local data samples’ size, it is necessary to split the local DNN model of each IoT device at different cut layers, and then offload different service-side models to the edge server to achieve device-edge synergy learning. In addition, in device-edge synergy learning, the network and computation resources’ limitation of the edge server shared by IoT devices, it is also necessary to efficiently allocate the sharing resource to further reduce the DNN training latency. Here, we formulate a joint model offloading and resource allocation problem as a mixed integers non-linear programming problem. Its main goal is to reduce the total training latency while satisfying data leakage risk rate constraints. To reduce the training latency for $N$ IoT devices in a round $t$ while satisfying data leakage risk rate constraints, we define our optimization target and constrained conditions for $N$ IoT devices as follows:
@@ 4 Problem formulation
where constraint condition $C_{1}$ in Eq. 19 is the data leakage risk rate constraint of each IoT device $I_{n}$, in which $P^{risk}$ is the risk rate constrain. Constraint condition $C_{2}$ is the time fraction allocation vector to $N$ IoT devices on the uplink channel. Constraint condition $C_{3}$ is the time fraction allocation vector to $N$ IoT devices on the downlink channel. Constraint condition $C_{4}$ denotes that the sum of computing resources allocated to $N$ IoT devices cannot exceed the computing capacity of the edge server. Constraint condition $C_{5}$ guarantees the cut layer of each IoT device $I_{n}$ is an integer.
The formulation of $\mathcal{P}_{1}$ is a maxed-integer non-linear programming problem (MINLP) which is difficult to solve [25]. To solve the formulation of $\mathcal{P}_{1}$ in a centralized algorithm itself is a detriment on privacy. That is because that the centralized algorithms are generally built on the complete knowledge regarding all IoT devices’ multiple dimensional training configurations, including the computation capacity, the mini-batch size, the size of local dataset and the number of epochs, etc. Hence, it may not suitable for many real-world smart applications exploiting SplitFed Learning. It is urgent to design a novel DP-MORA scheme which enables each IoT device to decide its own cut layer and resource requirement according to its own multiple dimensional training configuration without knowing other IoT devices’ multiple dimensional training configurations.
5 Proposed DP-MORA optimization algorithm
To enable each IoT device to decide its own cut layer and resource requirement according to its own multiple dimensional training configuration, we propose a DP-MORA scheme based on the block coordinate descent (BCD) method [26]. We first relax the formulation of $\mathcal{P}_{1}$ into the formulation of $\mathcal{P}_{2}$, then demonstrate the designing of the DP-MORA scheme.
For $\forall I_{n}\in\mathcal{I}$, we relax the discrete variable into continuous variable $\hat{\alpha_{n}}$. The relaxed formulation of $\mathcal{P}_{1}$ is denoted as follows.
@@ 5 Proposed DP-MORA optimization algorithm
Let $\nabla y(x)$ denote the partial derivative of function $y$ coresponding to variable $x$. Let $\text{Proj}_{\mathcal{X}}(x)=\text{arg}\min_{\kappa\in\mathcal{X}}||x-\kappa||^{2}$ denote the Euclidean projection of $x$ onto $\mathcal{X}$. The procedure of our proposed solution is presented in detail in Algorithm 1, and can be summerized as:
(1) Given $\mu_{d,n}^{DL}$, $\mu_{d,n}^{UL}$ and $\theta_{d}^{n}$, the decision variable of the optimization problem $\mathcal{P}_{2}$ is $\hat{\alpha}_{n}$. Each IoT device $I_{n}$ can derive its new $\hat{a}_{n}$ according to
@@
where $\eta_{\hat{a}_{n}}>0$ is a constant step size and ${\mathcal{X}_{\hat{a}_{n}}}$ is the bounded domain constrained by $\hat{C}_{5}$ of Problem $\mathcal{P}_{2}$. Based on the BCD method, we repeat Eq. 21 until the derived $\hat{a}_{n}$ is converged and then update $\hat{a}_{n}$. ($\hat{a}_{n}$ can converge to a local optimal solution)
@@
where constraints $\hat{C}_{1},C_{3},C_{4},\hat{C}_{5}$ are irrelevant to this problem. Due to the time fraction allocation constraint $C_{2}$ on the downlink channel shared by $N$ IoT devices, these $N$ IoT devices are coupled with each other, which leads to the method for solving $\hat{\alpha}_{n}$ not applied for solving $\mu_{d,n}^{DL}$. To address the optimization problem $\mathcal{P}_{2}$, a decentralized proactive downlink bandwidth resource allocation scheme is designed, the procedure of which is presented in detail in Algorithm 2. In order to eliminate the coupling relationship of $N$ IoT devices incurred by constraint $C_{2}$ and make $N$ IoT devices independently make decision its $\mu_{d,n}^{DL}$, the constraint $C_{2}$ can be further rewritten as
@@
where the constraint $C_{1}$ is to guarantee $\lambda_{1}=\ldots=\lambda_{n}=\ldots=\lambda_{N}$, which represents that the consensus is reached. Based on above reformulation, the optimization problem $P_{3}$ can be decoupled and each IoT device can decide its own time fraction allocations on the downlink channel of edge server.
@@
where $\mathcal{A}_{n}$ denotes the set of other IoT devices connected to IoT device $I_{n}$. Hence, to reach the global solution, each IoT device $I_{n}$ needs to know the information on $\lambda_{m}$, $z_{m}$, $I_{m}\in\mathcal{A}_{n}$. In our scenario, each IoT device cannot directly communicate with each other. However, since they are all connected to the edge server, they can utilize the edge server to relay the information on $\lambda_{m}$, $z_{m}$, $I_{m}\in\mathcal{A}_{n}$ to IoT device $I_{n}$. Based on the gradients $\nabla\mu_{d,n}^{DL}$, $\nabla\lambda_{n}$ and $\nabla z_{n}$, $\mu_{d,n}^{DL}$, $\lambda_{n}$ and $z_{n}$ can be updated by
@@
where $\eta$ represents integration step.
(3) Given the latest value $\hat{a}_{n}$ updated in step (1), the latest value $\mu_{d,n}^{DL}$ updated in step (2) and $\theta_{d}^{n}$, the decision variable of the optimization problem $\mathcal{P}_{2}$ is $\mu_{d,n}^{UL}$. Due to the time fraction allocation constraint $C_{3}$ on the uplink channel shared by $N$ IoT devices, these IoT devices are also coupled with each other. In addition, the problem formulation about the decision variable $\mu_{d,n}^{UL}$ is the same to that about the decision variable $\mu_{d,n}^{DL}$ in step (2). Hence, we adopt the same decentralized and privacy-preserving uplink bandwidth resource allocation scheme to solve the decision variable $\mu_{d,n}^{DL}$.
(4) Given the latest value $\hat{a}_{n}$ updated in step (1), the latest value $\mu_{d,n}^{DL}$ updated in step (2) and the latest value $\mu_{d,n}^{DL}$ updated in step (3), the decision variable of the optimization problem $\mathcal{P}_{2}$ is $\theta_{d}^{n}$. Similarity, due to the constraint $C_{4}$ on the edge server’s computing capacity shared by $N$ IoT devices, these IoT devices are also coupled with each other. Moreover, the problem formulation about the decision variable $\theta_{d}^{n}$ is the same to that about the decision variable $\mu_{d,n}^{DL}$ in step (2). Hence, we adopt the same decentralized and privacy-preserving computation resource allocation scheme to solve the decision variable $\theta_{d}^{n}$.
@@
4:     The edge server broadcasts the $\lambda_{n}(j)$, $z_{n}(j),\forall I_{n}\in\mathcal{I}$ to other IoT devices, $\mathcal{A}_{n}$;
5:     These $N$ IoT devices update and get $\mu_{d,n}^{DL}(j+1)$, $\lambda_{n}(j+1)$,$z_{n}(t+1)$;
6:     These $N$ IoT devices transfer $\lambda_{n}(j+1)$, $z_{n}(j+1)$ to the edge server; $j=j+1;$
@@
6 Theoretical analysis
6.1 The convergence and optimality of subproblems
@@
  • The information exchange among IoT devices can be formulated as a indirected and connected graph.
Proof. Proposition 1 gives three conditions for Algorithm 2 converging to the global optimal solution with any initial condition. If we can prove that these conditions hold for Algorithm 2, then we prove that Algorithm 2 can convergence to the global optimal solution with any initial condition. In the following, we prove that these three condition hold.
@@
  • Since $N$ IoT devices are all connected to the edge server and the edge server can relay the information among IoT devices, the information exchange among these IoT devices can be modeled as undirected and connected graph with assistant of the edge server. Hence, the third condition in proposition 1 holds for Algorithm 1.
So far, we have proved that these three conditions in proposition 1 hold. Thus, we have proved that Algorithm 2 can convergence to the global optimal solution with any initial condition.
6.2 The convergence of global problem
Proposition 2. Algorithm 2 based on the BCD method is proposed to solve the global optimization problem including four subproblems. According to [26], Algorithm 2 based on the BCD method can converges if at least two of subproblems are strictly quasi-convex.
Proof. Suppose that there exist $x_{1}$, $x_{2}$ and $\lambda$, where $\lambda\in(0,1)$, such that the strictly convex function is not strictly quasi convex function, we can derive
@@ 6.2 The convergence of global problem
We can find that Eq. 38 contradicts with the definition of strictly convex, where $f(\lambda x_{1}+(1-\lambda)x_{2})<\lambda f(x_{1})+(1-\lambda)f(x_{2})$, such that our above assumption is not true. As a result, there are three strictly quasi-convex subproblems. Hence, the condition for Algorithm 2 converging is satisfied. We thus complete the proof. Moreover, the subproblem respect with $\alpha_{n}$ is not convex. Therefore, the global problem does not necessarily converge to a global optimal solution, but a local optimal solution.
In section 6, we have proved that the optimization problem $\mathcal{P}_{1}$ is strongly convex with respect to $\mu_{d,n}^{DL}$, $\mu_{d,n}^{UL}$ and $f_{s}^{n}$. We adopt the contraction to prove strictly convex function be strictly quasi convex function.
7 Performance evaluation
In this section, we conduct extensive experiments to evaluate the performance of the proposed DP-MORA scheme. We first introduce the related experimental settings including IoT-edge environment parameters, training datasets, DNN models and several baseline algorithms. Then, we verify the efficiency and accuracy of the proposed DP-MORA scheme. Finally, we compare the DP-MORA scheme with these baseline approaches under different experimental settings, and analyze their impacts on the DP-MORA scheme.
7.1 Experimental settings
IoT-Edge Environment: The IoT-edge environment is consisting of an edge server and 10 IoT devices. By default, the computing capacity of the edge server is set to 60GFLOPS. The radio spectrum bandwidths for the edge server’s downlink and uplink are set to 50Mbps and 100Mbps, respectively. As IoT devices’ resource capacities are heterogeneous, we set three types of IoT devices: (1) Raspberry Pi3 with 3.62GFLOPS CPU-cycle frequency and 4GB memory in total, denoted as rpi3; (2) Raspberry Pi-3A+ with 5.0GFLOPS CPU-cycle frequency and 4GB memory in total, denoted as rpi3A+; (3) Raspberry Pi-4B (4GB) with 9.69GFLOPS CPU-cycle frequency and 4GB memory in total, denoted as rpi4B. These 10 IoT devices is consisting of 4 rpi3s, rpi3A+s and rpi4Bs. These 10 IoT devices can communicate with the edge server over the wireless network.
@@ 7.1 Experimental settings
DNN Models: In IoT-edge environment, the edge server cooperates with these heterogeneous IoT devices to train a global DNN model. Referring to the literature [16], we select RestNet 18 [30] and ResNet 34 two models as the evaluation models. That is because that the sample data can be recovered partially based on the gradients of server-side model [16]. The RestNet 18 is consisting of a convolution (CONV) layer, a max-pooling (POOL) layer, eight BasicBlocks and a fully-connected (FC) layers. The first two, the second two, the third two and the fourth two of these eight BasicBlocks are two $3*3*64$, two $3*3*128$, two $3*3*256$, and two $3*3*512$ convolution operations, respectively. The ResNet 34 is consisting of a convolution (CONV) layer, a max-pooling (POOL) layer, sixteen BasicBlocks and a fully-connected (FC) layers. The first three, the next four, the next six and the next three of these sixteen BasicBlocks are three $3*3*64$, four $3*3*128$, six $3*3*256$, and three $3*3*512$ convolution operations, respectively.
@@
Benchmarks: We first select four typical split Federated Learning scheme:
@@
(1) FedAvg [31]: The full model on each IoT device is trained on in parallel and their model parameters are aggregated to the edge server to obtain the updated global model.
@@
(2) SplitFed1 [8]: The full model on each IoT device is splited at the same cut layer. The identical server-side sub-model is offloaded to the edge server from each IoT device for device-edge synergy training in a sequential way.
@@
(3) SplitFed2: It adopts the same model offloading strategy as us, and multiple IoT devices train different sub-models in a sequential manner.
@@
(4) FederSplit [14]: Each IoT device offloads the identical server-side sub-model to the edge server for device-edge synergy training in a parallel manner.
@@
(5)SplitFed3: It adopts the same model offloading strategy as us, but multiple IoT devices train different sub-models in a parallel manner.
@@
Existing researches mainly focus on identifying the optimal cut layers for IoT devices. They fail to consider the impact of the edge server’s computation and network resource allocation on the training efficiency. In this paper, we jointly consider the problem of cut layer selection and resource allocation. Since there is no related resource allocation policy, we select two typical resource allocation scheme:
@@
(1) AF: It stands for the average resource allocation, which equally allocates the computing resource of edge server and the communication resource to $N$ IoT devices.
@@
(2) PF: It is the abbreviation of proportional fair, which allocates the computing resource of edge server and the communication resource according to the mini-batch sizes of $N$ IoT devices.
@@
Finally, we combine them to obtain eight joint offloading and resource allocation strategies: FedAvg+AF (FAAF), SplitFed1+PF (SF1PF), SplitFed1+AF (SF1AF), SplitFed2+PF (SF2PF), SplitFed2+AF (SF2AF), FederSplit+PF (FSPF), FederSplit+AF (FSAF), SplitFed3+PF(SF3PF), SplitFed3+AF(SF3AF). To demonstrate the proposed DP-MORA scheme superiority on training efficiency while satisfying the data leakage risk rate constraint, we compare the DP-MORA scheme with the above these baseline algorithms.
7.2 Performance evaluation of the proposed DP-MORA scheme

### v2

@@ III-D Regression-based modeling methodology
IV Problem formulation
In a resource-constrained heterogeneous IoT-edge computing environment, considering the heterogeneity of end devices’ resource capacities and their local data samples’ size, it is necessary to split the local DNN model of each end device at different cut layers, and then offload different server-side models to the edge server to achieve device-edge synergy learning. In device-edge synergy learning, due to the limited network and computation resources of edge server, it is also necessary to efficiently allocate the sharing resource to further reduce the DNN training latency. Here, we formulate the joint model offloading and resource allocation problem to be the problem $P_{1}$. Its main goal is to reduce the total training latency for $N$ end devices in a round $t$ while satisfying data leakage risk rate constraints. We define our optimization objective and constrained conditions as follows:
@@ IV Problem formulation
where $\tau_{n}^{t}(l_{n},\mu_{d,n}^{DL},\mu_{d,n}^{UL},\theta_{d}^{n})$ is the overall training latency of end device $I_{n}$ in a round $t$. Constraint condition $C_{1}$ in Eq. 19 is the data leakage risk rate constraint of each end device $I_{n}$, in which $P^{risk}$ is the risk rate constraint. Constraint condition $C_{2}$ is the time fraction allocation vector to $N$ end devices on the uplink channel. Constraint condition $C_{3}$ is the time fraction allocation vector to $N$ end devices on the downlink channel. Constraint condition $C_{4}$ denotes that the sum of computing resources allocated to $N$ end devices cannot exceed the computing capacity of the edge server. Constraint condition $C_{5}$ guarantees the cut layer of each end device $I_{n}$ is an integer.
The formulation of $\mathcal{P}_{1}$ is a maxed-integer non-linear programming problem (MINLP) which is difficult to solve [25]. To solve the formulation of $\mathcal{P}_{1}$ in a centralized algorithm itself is a detriment on privacy. That is because the centralized algorithms are generally built on the complete knowledge regarding all end devices’ multiple dimensional training configurations, including the computation capacity, the mini-batch size, the size of the local dataset and the number of epochs, etc. Hence, it may not be suitable for many real-world smart applications exploiting SplitFed Learning. It is urgent to design a novel DP-MORA scheme which enables each end device to decide its own cut layer and resource requirement according to its own multiple-dimensional training configuration without knowing other end devices’ multiple-dimensional training configurations.
V Proposed DP-MORA optimization algorithm
The block coordinate descent (BCD) method is to divide the variables of the optimization problem into multiple blocks, and only optimize the variables of one block in each iteration, while keeping the other variables fixed. Gradually optimizing a block of variables can result in much less computation than directly optimizing all variables. Therefore, BCD method is suitable to solve complex optimization problems with different properties about different variables (i.e., the convex problem with some variables or non-convex problem with another variable). Moreover, BCD shows good convergence in practice. In convex optimization problems, BCD is usually able to guarantee global convergence, while in non-convex problems, BCD is able to converge to a local optimal solution. In many practical problems, BCD can guarantee to converge to a local optimal solution by gradually optimizing the variable blocks. The local optimization of BCD on each subproblem makes the objective function value gradually decrease. Although the global optimum is not guaranteed, a better solution is usually obtained. Our formulated problem $P_{1}$ is a complex optimization problem with multiple variables. Subproblem with respect to different variables have different properties. For example, subproblem with variable $\mu_{d,n}^{DL}$ is convex problem, while the subproblem with respect to the variable $l_{n}$ is non-convex problem. The BCD algorithm has achieved excellent performance in solving this kind of problem. To achieve this, we propose a DP-MORA scheme based on the block coordinate descent (BCD) method [26], which enables each end device to decide its own cut layer and resource requirement according to its own multiple-dimensional training configurations. We first relax the formulation of $\mathcal{P}_{1}$ into the formulation of $\mathcal{P}_{2}$, then demonstrate the designing of the DP-MORA scheme.
$\forall I_{n}\in\mathcal{I}$, we relax the discrete variable ${\alpha}_{n}$ into continuous variable $\hat{\alpha}_{n}$. The relaxed formulation of $\mathcal{P}_{1}$ is denoted as follows:
@@ V Proposed DP-MORA optimization algorithm
Let $\nabla y(x)$ denote the partial derivative of function $y$ corresponding to variable $x$. Let $\text{Proj}_{\mathcal{X}}(x)=\text{arg}\min_{\kappa\in\mathcal{X}}||x-\kappa||^{2}$ denote the Euclidean projection of $x$ onto $\mathcal{X}$. The procedure of our proposed solution is presented in detail in Algorithm 1, and can be summarized as:
(1) Given $\mu_{d,n}^{DL}$, $\mu_{d,n}^{UL}$ and $\theta_{d}^{n}$, the decision variable of the optimization problem $\mathcal{P}_{2}$ is $\hat{\alpha}_{n}$. Each end device $I_{n}$ can derive its new $\hat{a}_{n}$ according to
@@
where $\eta_{\hat{a}_{n}}>0$ is a constant step size and ${\mathcal{X}_{\hat{a}_{n}}}$ is the bounded domain constrained by $\hat{C}_{5}$ of Problem $\mathcal{P}_{2}$. Based on the BCD method, we repeat Eq. 21 until the derived $\hat{a}_{n}$ is converged and then update $\hat{a}_{n}$. ($\hat{a}_{n}$ can converge to a local optimal solution).
@@
where constraints $\hat{C}_{1},C_{3},C_{4},\hat{C}_{5}$ are irrelevant to this problem. Due to the time fraction allocation constraint $C_{2}$ on the downlink channel shared by $N$ end devices, these $N$ end devices are coupled with each other, which leads to the method for solving $\hat{\alpha}_{n}$ not applied for solving $\mu_{d,n}^{DL}$. To address the optimization problem $\mathcal{P}_{2}$, a decentralized proactive downlink bandwidth resource allocation scheme is designed, the procedure of which is presented in detail in Algorithm 2. In order to eliminate the coupling relationship of $N$ end devices incurred by constraint $C_{2}$ and make $N$ end devices independently make decision its $\mu_{d,n}^{DL}$, the constraint $C_{2}$ can be further rewritten as
@@
where the constraint $C_{1}$ is to guarantee $\lambda_{1}=\ldots=\lambda_{n}=\ldots=\lambda_{N}$, which represents that the consensus is reached. Based on the above reformulations, the optimization problem $P_{3}$ can be decoupled and each end device can decide its own time fraction allocations on the downlink channel of edge server.
@@
where $\mathcal{A}_{n}$ denotes the set of other end devices connected to end device $I_{n}$. Hence, to reach the global solution, each end device $I_{n}$ needs to know the information on $\lambda_{m}$, $z_{m}$, $I_{m}\in\mathcal{A}_{n}$. In our scenario, each end device cannot directly communicate with each other. However, since they are all connected to the edge server, they can utilize the edge server to relay the information on $\lambda_{m}$, $z_{m}$, $I_{m}\in\mathcal{A}_{n}$ to end device $I_{n}$. Based on the gradients $\nabla\mu_{d,n}^{DL}$, $\nabla\lambda_{n}$ and $\nabla z_{n}$, $\mu_{d,n}^{DL}$, $\lambda_{n}$ and $z_{n}$ can be updated by
@@
where $\eta$ represents the integration step.
(3) Given the latest value $\hat{a}_{n}$ updated in step (1), the latest value $\mu_{d,n}^{DL}$ updated in step (2) and $\theta_{d}^{n}$, the decision variable of the optimization problem $\mathcal{P}_{2}$ is $\mu_{d,n}^{UL}$. Due to the time fraction allocation constraint $C_{3}$ on the uplink channel shared by $N$ end devices, these end devices are also coupled with each other. In addition, the problem formulation about the decision variable $\mu_{d,n}^{UL}$ is the same to that about the decision variable $\mu_{d,n}^{DL}$ in step (2). Hence, we adopt the same decentralized and privacy-preserving uplink bandwidth resource allocation scheme to solve the decision variable $\mu_{d,n}^{DL}$.
(4) Given the latest value $\hat{a}_{n}$ updated in step (1), the latest value $\mu_{d,n}^{DL}$ updated in step (2) and the latest value $\mu_{d,n}^{DL}$ updated in step (3), the decision variable of the optimization problem $\mathcal{P}_{2}$ is $\theta_{d}^{n}$. Similarly, due to the constraint $C_{4}$ on the edge server’s computing capacity shared by $N$ end devices, these end devices are also coupled with each other. Moreover, the problem formulation about the decision variable $\theta_{d}^{n}$ is the same to that about the decision variable $\mu_{d,n}^{DL}$ in step (2). Hence, we adopt the same decentralized and privacy-preserving computation resource allocation scheme to solve the decision variable $\theta_{d}^{n}$.
@@
4:     The edge server broadcasts the $\lambda_{n}(j)$, $z_{n}(j),\forall I_{n}\in\mathcal{I}$ to other end devices, $\mathcal{A}_{n}$;
5:     These $N$ end devices update and get $\mu_{d,n}^{DL}(j+1)$, $\lambda_{n}(j+1)$,$z_{n}(t+1)$;
6:     These $N$ end devices transfer $\lambda_{n}(j+1)$, $z_{n}(j+1)$ to the edge server; $j=j+1;$
@@
VI Theoretical analysis
VI-A The convergence and optimality of subproblems
@@
  • The information exchange among end devices can be formulated as an undirected and connected graph.
Proof. Proposition 1 gives three conditions for Algorithm 2 to converge to the global optimal solution with any initial condition. If we can prove that these conditions hold for Algorithm 2, then we prove that Algorithm 2 can converge to the global optimal solution with any initial condition. In the following, we prove that these three conditions hold.
@@
  • Since $N$ end devices are all connected to the edge server and the edge server can relay the information among end devices, the information exchange among these end devices can be modeled as an undirected and connected graph with assistance of the edge server. Hence, the third condition in proposition 1 holds for Algorithm 1.
So far, we have proved that these three conditions in proposition 1 hold. Thus, we have proved that Algorithm 2 can converge to the global optimal solution with any initial condition.
VI-B The convergence of global problem
Proposition 2. Algorithm 2 based on the BCD method is proposed to solve the global optimization problem including four subproblems. According to [26], Algorithm 2 based on the BCD method can converge if at least two of the subproblems are strictly quasi-convex.
Proof. Suppose that there exists $x_{1}$, $x_{2}$ and $\lambda$, where $\lambda\in(0,1)$, such that the strictly convex function is not strictly quasi convex function, we can derive
@@ VI-B The convergence of global problem
We can find that Eq. 38 contradicts the definition of strictly convex, where $f(\lambda x_{1}+(1-\lambda)x_{2})<\lambda f(x_{1})+(1-\lambda)f(x_{2})$, such that our above assumption is not true. As a result, there are three strictly quasi-convex subproblems. Hence, the condition for Algorithm 2 converging is satisfied. We thus complete the proof. Moreover, the subproblem respect with $\alpha_{n}$ is not convex. Therefore, the global problem does not necessarily converge to a global optimal solution, but a local optimal solution.
In section 6, we have proved that the optimization problem $\mathcal{P}_{1}$ is strongly convex with respect to $\mu_{d,n}^{DL}$, $\mu_{d,n}^{UL}$ and $f_{s}^{n}$. We prove strictly convex function to be strictly quasi convex function by adopting the contradiction.
VII Performance evaluation
In this section, we conduct extensive experiments to evaluate the performance of the proposed DP-MORA scheme. We first introduce the related experimental settings including IoT-edge computing environment parameters, training datasets, DNN models and several baseline algorithms. Then, we verify the efficiency and accuracy of the proposed DP-MORA scheme. Finally, we compare the DP-MORA scheme with these baseline approaches under different experimental settings, and analyze their impacts on the DP-MORA scheme.
VII-A Experimental settings
IoT-Edge Computing Environment: The IoT-edge computing environment is consisting of an edge server and 10 end devices. By default, the computing capacity of the edge server is set to 60GFLOPS. The radio spectrum bandwidths for the edge server’s downlink and uplink are set to 50Mbps and 100Mbps, respectively. As end devices’ resource capacities are heterogeneous, we set three types of end devices: (1) Raspberry Pi3 with 3.62GFLOPS CPU-cycle frequency and 4GB memory in total, denoted as rpi3; (2) Raspberry Pi-3A+ with 5.0GFLOPS CPU-cycle frequency and 4GB memory in total, denoted as rpi3A+; (3) Raspberry Pi-4B (4GB) with 9.69GFLOPS CPU-cycle frequency and 4GB memory in total, denoted as rpi4B. These 10 end devices consist of 4 rpi3s, rpi3A+s and rpi4Bs. These 10 end devices can communicate with the edge server over the wireless network.
@@ VII-A Experimental settings
DNN Models: In an IoT-edge computing environment, the edge server cooperates with these heterogeneous end devices to train a global DNN model. Referring to the literature [16], we select ResNet 18 [30] and ResNet 34 two models as the evaluation models. That is because the sample data can be recovered partially based on the gradients of server-side model [16]. The ResNet 18 consists of a convolution (CONV) layer, a max-pooling (POOL) layer, eight BasicBlocks and a fully-connected (FC) layers. The first two, the second two, the third two and the fourth two of these eight BasicBlocks are two $3*3*64$, two $3*3*128$, two $3*3*256$, and two $3*3*512$ convolution operations, respectively. The ResNet 34 consists of a convolution (CONV) layer, a max-pooling (POOL) layer, sixteen BasicBlocks and a fully-connected (FC) layers. The first three, the next four, the next six and the next three of these sixteen BasicBlocks are three $3*3*64$, four $3*3*128$, six $3*3*256$, and three $3*3*512$ convolution operations, respectively.
@@
Benchmarks: We first select four typical split Federated Learning schemes:
@@
(1) FedAvg [31]: The full model on each end device is trained in parallel and their model parameters are aggregated to the edge server to obtain the updated global model.
@@
(2) SplitFed1 [8]: The full model on each end device is splited at the same cut layer. The identical server-side sub-model is offloaded to the edge server from each end device for device-edge synergy training in a sequential way.
@@
(3) SplitFed2: It adopts the same model offloading strategy as ours, and multiple end devices train different sub-models in a sequential manner.
@@
(4) FederSplit [14]: Each end device offloads the identical server-side sub-model to the edge server for device-edge synergy training in a parallel manner.
@@
(5)SplitFed3: It adopts the same model offloading strategy as ours, but multiple end devices train different sub-models in a parallel manner.
@@
Existing researches mainly focus on identifying the optimal cut layers for end devices. They fail to consider the impact of the edge server’s computation and network resource allocation on the training efficiency. In this paper, we jointly consider the problem of cut layer selection and resource allocation. Since there is no related resource allocation policy, we select two typical resource allocation schemes:
@@
(1) Average Fair (AF): It stands for the average resource allocation, which equally allocates the computing resource of the edge server and the communication resource to $N$ end devices.
@@
(2) Proportional Fair (PF): It is the abbreviation of proportional fair, which allocates the computing resource of the edge server and the communication resource according to the mini-batch sizes of $N$ end devices.
@@
Finally, we combine them to obtain eight joint offloading and resource allocation strategies: FedAvg+AF(FAAF), SplitFed1+PF (SF1PF), SplitFed1+AF(SF1AF), SplitFed2+PF(SF2PF), SplitFed2+AF (SF2AF), FederSplit+PF(FSPF), FederSplit+AF(FSAF), SplitFed3+PF(SF3PF), SplitFed3+AF(SF3AF). To demonstrate the proposed DP-MORA scheme’s superiority in training efficiency while satisfying the data leakage risk rate constraint, we compare the DP-MORA scheme with these above baseline algorithms.
VII-B Performance evaluation of the proposed DP-MORA scheme